Solve the given equation, check your solution.
-3|2n+5|=-9

Answer by jim_thompson5910(35256)

-3|2n+5|=-9

|2n+5|=-9/(-3)

|2n+5| = 3

2n+5 = 3 or 2n+5 = -3

2n = 3-5 or 2n = -3-5

2n = -2 or 2n = -8

n = -2/2 or n = -8/2

n = -1 or n = -4

The two solutions are n = -1 or n = -4